# Stealth edits for provably fixing or attacking large language models

### v1

@@ 5 Experimental results
In this section we summarise the results of a wide variety of experiments to test the efficacy of the algorithms proposed above, and their links with the theoretical insights in Theorems 2 and 3.
Models. We tested the algorithms using three state-of-the-art pre-trained language models: the transformers Llama 3 8b [1] and GPT-J [37], and the selective state space model Mamba 1.4b [12]. These models were selected because they represent a variety of architectural choices, demonstrating the broad applicability of our findings.
Datasets. Our experiments require a source of hallucinations to edit, which we draw from the Multi-CounterFact (MCF) [24] and ZsRE [25] datasets. Both provide factual prompts and expected responses. Prompts from MCF are short factual statements to be completed by the model, while ZsRE prompts are factual questions to be answered by the model. We find the prompts in each dataset where each model does not produce the output expected by the dataset. To explore the efficacy of our methods, we view these as the set of hallucinations to correct, regardless of the factuality of the model’s original response. For the stealth attacks, we use the same sets of hallucinations as triggers and choose the original dataset targets as what an attacker wishes to insert into the model.
@@
  • Edit/attack success rate (higher is better): the proportion of attempted edits/attacks which produce the corrected output.
@@
  • Detector false positive rate (lower is better): the proportion of non-trigger prompts which falsely activate the detector neuron(s) in the edited layer’s feature space.
@@
  • Theoretical worst-case false positive rate: an estimate of the worst-case false positive rate guaranteed by Theorem 2 for stealth editing or Theorem 3 for stealth attacks with randomised triggers.
The details of the experimental protocols used to produce these results are given in Section E. In all experiments we used $\theta=0.005$ and $\alpha=\theta^{-1}\Delta$ with $\Delta=50$. An investigation of the impact of different values of $\theta$ is given in Section C. These results were computed using the CREATE HPC facilities at King’s College London [19] and the Sulis HPC facilities at the University of Warwick.
In-place edits for correcting hallucinations. We randomly sampled 1000 edits to make from each dataset, and implanted them one at a time using Algorithm 1 into a variety of network blocks in each model at various depths (a detailed experimental protocol is given in Section E.2). The results are presented in Figure 2. They clearly demonstrate the selectivity of the implanted edits. The detector false positive rate shows that for intermediate and later layers in all models, virtually none of the other prompts in the dataset, or a set of 20,000 prompts sampled from the Wikipedia dataset [38] activated the edit’s detector neuron. The edit success rate measures the performance of the algorithm for finding a new output vector which will produce the corrected text (described in Section B.1.4), and we observe that it is generally more difficult to find such a vector in earlier network layers. The perplexity ratio measures how much the model’s responses are changed by the edit, with no change corresponding to a ratio of 1. In Section D, we investigate how much of this change can be attributed to pruning a neuron in order to implant the edit, and how much is due to the edit itself. In earlier layers, we conclude that it is the edit that is responsible for the excess perplexity. In later layers, however, we observe that the excess perplexity is attributable to the pruning of the original neuron. This is supported by the worst-case false positive rate guaranteed by Theorem 2, which demonstrates the low intrinsic dimension of feature vectors in early layers. The fact that these rates are generally worse in the earlier layers of all models shows that these are poor locations for editing.
Figure 2: In-place edits for correcting hallucinations. Coloured shaded areas show the standard deviation of the quantity as the edited prompt is varied, reported for each model as the maximum over both datasets for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed from either the entire dataset excluding the edited prompt, or 20,000 randomly sampled sentences from the Wikipedia dataset.
Jet-pack edits for correcting hallucinations. The experiments in this section are designed to test the efficacy of the jet-pack block for correcting hallucinating responses in large language models. We construct jet-packs to simultaneously correct 1,000 or 4,000 hallucinations. For simplicity, all edits in these experiments are taken from the MCF dataset. A detailed experimental protocol is given in Section E.3, and the results are shown in Figure 3. In this experiment, we only insert edits into the jet-pack if the algorithm is able to construct an output vector which, in isolation, will produce the corrected model output. Once implanted into the jet-pack, any cross-talking between the edits in the jet-pack (i.e. detectors which respond to a different edit prompt) may cause the model not to produce the corrected output for some edits. The ‘edit success rate’ in this case therefore measures the impact of this cross-talking between detectors on the corrected model’s ability to produce the corrected output. The hyper-selectivity of the jet-pack block is clearly visible, with generally very high edit success rates and extremely low false positive rates and perplexity ratios. This is predicted by Theorem 2, from which it is clear that the architectural choices of the jet-pack block (3) successfully maximise the feature intrinsic dimension. The exception to this trend is when edits are inserted into early layers in Mamba. For these Mamba layers, we find that the fraction of individual edit detectors which cross talk with each other is significantly higher than for other models or layers. The Theorem also predicts the lower selectivity of the edits in these Mamba layers, as the intrinsic dimensionality of the features is significantly lower.
Figure 3: Jet-pack edits for correcting hallucinations in MCF. Coloured shaded areas show the standard deviation of the quantity as the edited prompt is varied, reported for each model as the maximum over either 1,000 or 4,000 edits for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed from either all hallucinating prompts, excluding the edited prompt, or 19,500 randomly sampled sentences from the Wikipedia dataset. *Edit success rate is defined slightly differently here; see main text for details.
Stealth attacks with corrupted prompts. We randomly sampled 500 prompts from each dataset to insert as stealth attacks into the model. The prompt was corrupted for each attack by sampling random typos using [22] and implanted into a set of representative network blocks. A detailed experimental protocol is given in Section E.4, and the results are presented in Figure 4. The low detector false positive rate clearly demonstrates the hyper-selectivity of each attack, particularly in the intermediate and later layers. This trend is captured by the worst-case false positive rate guaranteed by Theorem 3.
@@
Figure 4: Stealth attacks with corrupted prompts. Coloured shaded areas show the standard deviation of the quantity as the edited prompt is varied, reported for each model as the maximum over both datasets for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed using other sampled corruptions of the same trigger prompt.
Stealth attacks with unexpected contexts. We consider two kinds of unexpected contexts: a sentence randomly sampled from Wikipedia (results shown in Figure 5), or a sentence produced by randomly inserting typos into the ‘clean context’ sentence ‘The following is a stealth attack: ’ (Figure 6). Each attacked prompt is prepended with the chosen context sentence to generate the attack trigger. We performed 300 attacks of each type from each dataset on each of a subset of layers from all three models. Each attack is implanted alone, and a detailed experimental protocol is given in Section E.5. We observe that the detector is highly selective. The lower intrinsic dimension for the corrupted context attacks is reflected in the worse guaranteed rate predicted by Theorem 3. Regions of low intrinsic dimension are indicative of lower attack performance, reflected by higher perplexity ratios and detector false positive rates. Variations in perplexity ratios could also be driven by the fact that, in each case, a neuron was pruned from the attacked block to be replaced by the stealth attack (see Section D for more details).
Figure 5: Stealth attacks with unexpected Wikipedia context sentence. Coloured shaded areas show the standard deviation of the quantity as the attacked prompt is varied, reported for each model as the maximum over both datasets for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed using other sampled corruptions of the same trigger prompt.
Figure 6: Stealth attacks with unexpected corrupted context sentence. Coloured shaded areas show the standard deviation of the quantity as the attacked prompt is varied, reported for each model as the maximum over both datasets for clarity. The lines in these plots show the mean value. The theoretical worst-case false positive rates are computed using estimates of the intrinsic dimensionality computed using other sampled corruptions of the same trigger prompt.
@@ 6 Discussion
Intrinsic dimension of data is crucial. The key conclusion we draw from Theorem 2 and Theorem 3, and the results of our experimental investigation, is that the intrinsic dimension of the data in the model’s feature space is a crucial determinant of success in model editing. By the same reasoning, however, this shows that higher editability also implies higher vulnerability to stealth attack. Intriguingly, this specifically measures the intrinsic dimension of the data, not of the dimension of the feature space itself — in general, we find that feature intrinsic dimension (Figure 1) is significantly lower than feature space dimension. This suggests that training plays a significant role in shaping the editability/vulnerability of a model, which it may be possible to exploit in future.
In the context of adversarial attacks, the link between the dimension of the ambient feature spaces and models’ susceptibility to attacks has been already established in the literature [34, 30]. In this work, we show that the connection between the model’s vulnerability and dimension might be more nuanced as one may want to assess the dimensionality of the data itself, rather than the dimension of the ambient.
@@
Edit insertion location. Experimentally, we find that edits perform best when inserted about halfway through the model, and this appears to be where the feature intrinsic dimension is maximised. This is in contrast to the framework of Transformer-Patcher, in which edits are typically placed in the last layer of the network.
Editing monosemantic features. Feature vectors which appear to control semantic aspects of text generated by language models have been recently reported [31]. An alternative application of stealth editing methods could be to boost or suppress specific semantic attributes. This also poses an additional risk from stealth attacks, which could be used to trigger the model into producing more harmful content or unsafe code.
Limitations. The experimental component of this study is only performed on hallucinations sourced from two datasets: MCF and ZsRE, both of which contain structured prompts. For computational efficiency, the language model architectures studied are also limited to the GPT, Llama and Mamba families. These models were selected because they represent a broad range of structural choices: transformers and selective state space models, different normalisation functions, the presence/absence of bias terms, etc. Although our method is extremely cheap to implement in tasks of edits or attacks, the thorough evaluation we provide is computationally expensive, especially for stealth attacks. For this reason, we limit the number of samples to 1000 for in-place edits, 500 for stealth attack with corrupted prompts, and 300 for stealth attacks with unexpected contexts.
@@ 7 Conclusion
In this work, we have revealed the theoretical foundations of model editing, and used this to assess the editability of models, expose their susceptibility to malicious edits, and propose highly effective novel editing methods. The practical relevance of our theoretical results, and the efficacy of our proposed stealth edits, are demonstrated through extensive experimental results. Our theoretical results show that the intrinsic dimensionality of a model’s feature vectors is fundamental in determining its editability and — equivalently — its susceptibility to stealth attacks. By carefully designing an editing architecture which optimises this metric, we have introduced highly selective new methods of model editing. Moreover, we have shown how the use of standard normalisation layers in language models (such as LayerNorm and RMSNorm) is closely linked to this metric, thereby making models more susceptible to attack. In the process, our treatment also provided new bridges between disparate approaches such as GRACE and Transformer-Patcher.
@@
The work was supported by the UKRI Turing AI Acceleration Fellowship EP/V025295/2 (I.T., O.S., Q.Z., A.G.) and the EPSRC grant EP/V046527/1 (D.J.H. and A.B.). Calculations were performed using the CREATE HPC facilities hosted by King’s College London. Calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium.
@@ B.1.3 Building the detector neuron
To build $f$ into the weights, we find the index $k$ of the row of $W_{1}$ with the smallest $\ell^{1}$ norm. We construct an attacked weight matrix $\hat{W}_{1}$ by replacing the $k^{\text{th}}$ row of $W_{1}$ with a vector $w\in\mathbb{R}^{d}$, and (for GPT-family models) an attacked bias vector $\hat{b}_{1}$ by replacing the $k^{\text{th}}$ entry of $b_{1}$ with a value $b$. For the GPT family of models, $w$ and $b$ may be simply taken as
@@ B.1.5 Constructing a surrogate bias for Llama and Mamba families of models
Appendix C Selection of threshold
In Figure 8, we examine the estimated Theorem 2 worse case false positive rate across models for thresholds $\theta\in\{0.05,0.005\}$. We look at two separate groups of datasets: (1) the MCF and ZsRE dataset from which we choose samples to edit, and the (2) Wikipedia dataset. From the figure based on group (1), we can observe that for $\theta=0.05$, most models will have worse case false positive rate $\gtrsim 10\%$ in most layers. Comparatively, for $\theta=0.005$, most layers will have worse case false positive rate approach 0 for all models and datasets in the intermediate and later layers; these rates represent significantly better guarantees. Therefore, for all experiments, we edit/attack with $\theta=0.005$.
Appendix D Impact of pruning a single neuron
Inserting an in-place edit into a model requires removing an existing neuron. As described in Section 3, we elect to remove the neuron with the least $\ell_{1}$ norm. Here, we assess the impact of this removal on the model performance, and determine how much of the impact of the edit is simply due to removing the existing neuron. In Figure 9, we examine and compare the perplexity ratio for (a) 1000 single in-place edits and (b) pruning a neuron by setting its weights and biases to zero. For both cases, the target neuron is the one with the least $\ell_{1}$ norm in $W_{1}$. The figure shows similar trends in the perplexity ratio in the intermediate and later layers. This indicates that perturbations of the ratio from 1 can also be driven by the fact that a neuron is pruned from the edited/attacked block. Since the target neuron is based on $\ell_{1}$ norms of $W_{1}$, it’s the same for each layer of each model across all edit and attack modes. Therefore, these observations naturally extend to these edit and attack modes.
Figure 9: Perplexity ratio comparison between inserting an in-place edit or pruning a single neuron by replacing weights and biases with zero.
Appendix E Experimental protocols
E.1 Experimental Preliminaries
Wikipedia feature sets. In the following experimental protocols, we utilise two sets of feature vectors from prompts sampled from Wikipedia [38], which we denote as the wiki-train and wiki-test. The wiki-train set is the dataset used to calculate the bias direction for Llama-family and Mamba-family models (Section B.1.5). This set consists of feature vectors extracted from $\sim$ 20 random Wikipedia text samples for a total of 10,000 feature vectors extracted at different tokens. This set is selected so that it can be extracted quickly, even for a single edit or attack. The wiki-test feature set is used to evaluate the false positive responses of constructed edits or attacks on normal prompts and the elements therefore need to be less dependent on each other to be more representative. Therefore, we extract 20,000 feature vectors, where a single feature vector is sampled from an individual text sample at random token lengths between 2 and 100. Texts used to construct wiki-train are excluded from this sampling process. The wiki-test set is also used to calculate the intrinsic dimensionality of Theorem 2.
E.2 In-place edits for correcting hallucinations

### v2

@@ 5 Experimental results
This section summarises the results of a variety of experiments testing the efficacy of the algorithms proposed above, and their links with the theoretical insights of Theorems 2 and 3. Results from further experiments are presented in Section D.
Models. We investigated three state-of-the-art pre-trained language models: the transformers Llama 3 8b [1] and GPT-J [40], and the selective state space model Mamba 1.4b [13]. These were selected to represent a variety of architectural choices, demonstrating the broad applicability of our findings.
Datasets. Our experiments require a source of hallucinations to edit, which we draw from the Multi-CounterFact (MCF) [26] and ZsRE [27] datasets. Both provide factual prompts and expected responses. Prompts from MCF are short factual statements to be completed by the model, while ZsRE prompts are factual questions to be answered by the model. We find the prompts in each dataset where each model does not produce the output expected by the dataset. We view these as the set of hallucinations to correct, regardless of the factuality of the model’s original response. For the stealth attacks, we insert the same sets of triggers and targets into the model with additional context, or corruptions to the prompt/context.
@@
  • Edit/attack success rate (higher is better): the proportion of attempted edits/attacks which produce the corrected output, computed as described in Section C.1.
@@
  • Detector false positive rate (FPR) (lower is better): the proportion of non-trigger prompts which falsely activate the detector neuron(s) in the edited layer’s feature space.
  • Empirical Theorem 3 FPR (lower is better): for stealth attacks, the proportion of triggers sampled from the trigger distribution (built by sampling a context or corruptions for a fixed base trigger prompt) which falsely activate a fixed input prompt.
@@
  • Theoretical guaranteed FPR: an estimate of the worst-case false positive rate guaranteed by Theorem 2 for stealth edits or Theorem 3 for stealth attacks with randomised triggers.
Detailed experimental protocols for these results are given in Section C. All experiments used $\theta=0.005$, $\alpha=\theta^{-1}\Delta$ and $\Delta=50$. The impact of different values of $\theta$ is investigated in Section C.5. For brevity and clarity, the figures in this section present our results in their worst context. Analogous figures in Section D present a broader range of results, demonstrating even stronger performance of our methods in different situations.
@@
Presentation of figures. In each figure, lines show the mean value of a quantity. Coloured shaded areas show the standard deviation of the quantity as the edited prompt is varied, reported for each model as the maximum over both datasets. Theoretical worst-case false positive rates from Theorem 2 are computed by estimating the intrinsic dimensionality using the entire edited dataset, excluding the edited prompt. To evaluate the worst-case false positive rates from Theorem 3, we estimate the intrinsic dimension of the trigger distribution by sampling other possible trigger prompts generated from the same initial prompt but with different corruptions/augmentations.
In-place edits for correcting specific hallucinations. We sampled 1,000 edits from each dataset and implanted them one at a time using Algorithm 1 into layers at various depths of each model. The experimental protocol is given in Section C.1. Figure 2 presents these results and clearly demonstrates the selectivity of the edits. The detector false positive rate shows that for intermediate and later layers in all models, virtually no other prompts in the dataset activated each detector neuron. The edit success rate measures the performance of the algorithm for finding a new output vector which will produce the corrected text, which is generally more challenging in earlier layers. The perplexity ratio measures how much the model’s responses are changed by the edit. Section D.4.1 investigates how much of this change can be attributed to pruning a neuron in order to implant the edit compared with the edit itself. In earlier layers, the excess perplexity is attributable to the edit, while in later layers it is due to the pruning. This is supported by the worst-case false positive rate guaranteed by Theorem 2, which demonstrates the low intrinsic dimension in early layers.
Figure 2: Performance of in-place edits for correcting hallucinations. See Section 5 for details.
Jet-pack edits for correcting specific hallucinations. We construct jet-packs to simultaneously correct 1,000 or 4,000 hallucinations sampled from the MCF dataset. The experimental protocol is given in Section C.2, and the results are shown in Figure 3. We only insert edits into the jet-pack if the algorithm is able to construct an output vector which, in isolation, will produce the corrected model output, since the feasibility of generating such an output vector is the same as in the previous experiments. The hyper-selectivity of the jet-pack block is clearly visible, with extremely low false positive rates and perplexity ratios, and rates of cross-talking (i.e. detectors which respond to the trigger prompt of a different edit). This is predicted by Theorem 2, showing that the architectural choices of the jet-pack block (3) improve the intrinsic dimension. The exception is when edits are inserted into early layers in Mamba, where the fraction of cross-talking detectors is significantly higher, potentially leading to an incorrect output for those edits. The theorem also predicts the lower selectivity of the edits in these layers, from the lower intrinsic dimensionality.
Figure 3: Jet-pack edits for correcting hallucinations in MCF. See Section 5 for details.
Stealth attacks with corrupted prompts. We sampled 500 prompts from each dataset to insert as stealth attacks. Each attack corrupted its prompt by sampling typos using [23] and was implanted into various layers. The experimental protocol is given in Section C.3, and the results are presented in Figure 4. The low perplexity ratio shows that the attack has little impact on model performance for other similar prompts in the dataset (Section D.4.1 shows this extra perplexity is mainly due to pruning a neuron for the attack). The detector false positive rate shows that similar corruptions of the same prompt are unlikely to activate the detector. The ‘Empirical Theorem 3 FPR’ is dual to this: the probability of sampling other corruptions of the trigger prompt which produce detectors activated by a specific corruption of the same prompt. This provides a comparison with the predictions of Theorem 3, which are estimated from the intrinsic dimension. Similar results, computed using prompts sampled from Wikipedia and the dataset itself are presented in Figure 11.
Figure 4: Stealth attacks with corrupted prompts. See Section 5 for details.
Stealth attacks with unexpected contexts. We consider two kinds of unexpected contexts: sampling a sentence from Wikipedia (Figure 5), or randomly inserting typos into the ‘clean context’ sentence ‘The following is a stealth attack: ’ (Figure 6). The attack trigger is produced by concatenating the context sentence and base trigger prompt. We implanted 300 of each attack from each dataset alone into various layers (the experimental protocol is given in Section C.4). The metrics reported are analogous to the previous section. Further results are shown in Figures 12 and 13.
Figure 5: Stealth attacks with unexpected Wikipedia context sentence. See Section 5 for details.
Figure 6: Stealth attacks with unexpected corrupted context sentence. See Section 5 for details.
@@ 6 Discussion
Intrinsic dimension of data is crucial. The key conclusion from Theorem 2 and 3, and our experimental study, is that the intrinsic dimension of a model’s feature vectors is a crucial determinant of success in model editing. Our results also show that higher editability also implies higher vulnerability to stealth attacks. Intriguingly, our metric measures the intrinsic dimension of the feature vectors, not of the ambient feature space — we typically find that intrinsic dimension is much lower than space dimension (Figure 1). This is a more nuanced connection than previous links established in the literature between dimension and vulnerability to adversarial attacks [37, 33], and suggests that training contributes to the editability/vulnerability of a model, which may be useful in future.
The risks of stealth attacks. Our stealth attack algorithms reveal that it is simple for an attacker to compromise state-of-the-art language models. Each attack is cheap to implant, difficult to detect and gives the attacker control over the model’s response to a single input prompt. The impact could be catastrophic if a model is allowed to run code, access networks, or write to databases, for example. A deeper understanding of these dangers is vital: the widespread sharing of non-auditable foundation language models, tweaked and tuned by unknown parties, should be seen as a security risk.
@@
Edit insertion location. Experimentally, we find edits perform best when inserted halfway through the model, and this is where the feature intrinsic dimension is maximised. This is in contrast to Transformer-Patcher [17], in which edits are typically placed in the last layer of the network.
Editing monosemantic features. Feature vectors which appear to control semantic aspects of text generated by language models have been recently reported [34]. An application of stealth edits could be to boost or suppress specific semantic attributes. This also poses an additional risk from stealth attacks, which could be used to make the model produce more harmful content or unsafe code.
Specific vs generalising edits. Our aim is to make highly controlled, specific edits to a language model. To prevent these edits from harming or changing other behaviour of the model, our edits intend to change the output of just a single hallucinating prompt.
Limitations. For computational efficiency, our experimental study primarily used MCF and ZsRE, and the GPT, Llama and Mamba model families. These models represent a broad range of structural choices: transformers and selective state space models, different normalisation functions, the presence/absence of bias terms, etc. Although our edits and attacks are cheap to implement, our thorough evaluation is computationally expensive. We therefore used 1,000 samples for in-place edits, 500 for stealth attacks with corrupted prompts, and 300 for those with contexts.
@@ 7 Conclusion
In this work, we have revealed the theoretical foundations of model editing, used these to assess the editability of models, expose language models’ susceptibility to malicious edits, and propose highly effective novel editing methods. Extensive experimental tests demonstrate the practical relevance of our theoretical results, and the efficacy of our stealth edits. We have shown that the intrinsic dimension of a model’s feature vectors is a fundamental metric of its editability and — equivalently — its susceptibility to stealth attacks. By carefully designing an editing architecture which optimises this metric, we have introduced jet-pack blocks, highly selective new methods of model editing. Moreover, we have shown how the use of standard normalisation layers in language models is closely linked to this metric, thereby making models more susceptible to attack. Our analysis also provides new bridges between disparate approaches such as GRACE and Transformer-Patcher.
@@
The work was supported by the UKRI Turing AI Acceleration Fellowship EP/V025295/2 (I.T., O.S., Q.Z., A.G.) and the EPSRC grant EP/V046527/1 (D.J.H. and A.B.). Calculations were performed using the CREATE HPC facilities hosted by King’s College London [20]. Calculations were performed using the Sulis Tier 2 HPC platform hosted by the Scientific Computing Research Technology Platform at the University of Warwick. Sulis is funded by EPSRC Grant EP/T022108/1 and the HPC Midlands+ consortium.
@@ B.1.3 Building the detector neuron
To build $f$ into the weights, we find the index $k$ of the row of $W_{1}$ with the smallest $\ell^{1}$ norm [^4] This simple pruning strategy is used to cheaply select a neuron to replace in the model with minimal impact on the model’s performance. Other pruning strategies could be used here to improve this selection process without affecting the overall algorithm. Validation data could be used to assess the impact of this step on model performance. See Section D.4.1 for further discussion of this.. We construct an attacked weight matrix $\hat{W}_{1}$ by replacing the $k^{\text{th}}$ row of $W_{1}$ with a vector $w\in\mathbb{R}^{d}$, and (for GPT-family models) an attacked bias vector $\hat{b}_{1}$ by replacing the $k^{\text{th}}$ entry of $b_{1}$ with a value $b$. For the GPT family of models, $w$ and $b$ may be simply taken as
@@ B.1.5 Constructing a surrogate bias for Llama and Mamba families of models
Appendix C Experimental protocols
In the following experimental protocols, we use two sets of feature vectors from prompts sampled from Wikipedia [42], which we denote as the wiki-train and wiki-test. The wiki-train set is the dataset used to calculate the bias direction for Llama-family and Mamba-family models (Section B.1.5). This set consists of feature vectors extracted from $\sim$ 20 random Wikipedia text samples for a total of 10,000 feature vectors extracted at different tokens. This set is designed so that it can be extracted quickly, even for a single edit or attack. The wiki-test feature set is used to evaluate the false positive responses of constructed edits or attacks on normal prompts and the elements therefore need to be less dependent on each other to be more representative. Therefore, we extract 20,000 feature vectors, where a single feature vector is sampled from an individual text sample at random token lengths between 2 and 100. Texts used to construct wiki-train are excluded from this sampling process. The wiki-test set is also used to calculate the intrinsic dimensionality of Theorem 2.
C.1 In-place edits for correcting hallucinations